A human gene carries a certain disease from the mother to the child with a probability rate of 54%. That is, there is a 54% chance that the child becomes infected with the disease. Suppose a female carrier of the gene has five children. Assume that the infections of the five children are independent of one another. Find the probability that at least one of the children get the disease from their mother. Round to the nearest thousandth.
Answer #1

X ~ bin ( n , p)

Where n = 5 , p = 0.54

P(X) = nCx px ( 1 - p)n-x

P(X >= 1) = 1 - P(X = 0)

= 1 - 5C0 * 0.540 * ( 1 - 0.54)5

= 0.979